### v1

@@ 4.1 Part Localization evaluation
Table 2: Ablation Analysis of PartCraft for Part Localization.
PartCraft (Ablations) Method FG-NMI FG-ARI PartCraft 35.4 11.0 w/o Null-Text Inversion 23.1 5.2 w/o Max Localization 21.3 2.8 w/o Dot Product Localization 23.7 5.0 w/o Independent Text 31.2 8.9 PartCraft (Clustering) PartCraft (K = 9) 35.4 11.0 PartCraft (K = 4) 20.4 2.8 PartCraft (K = 14) 35.2 10.3
Figure 7: User Study and Quantitative Results for part-based image generation.
PartCraft StableDiffusion InstructPix2Pix Rich-Text LPIPS 0.168 0.467 0.189 0.243 CLIP 0.201 0.183 0.193 0.187 Aesthetic 5.66 5.68 5.63 5.65
Figure 8: PartCraft generalizes over domains as shown by Claude Monet’s painting image (above), where we specify the dress part of women to follow styles like Van Gogh and Ukiyo-e in different colors. We perform similar modifications to Woodie’s image as well, where it’s able to generate aesthetic novel compositions with specified parts.
4.2 PartCraft Image Generation Evaluation
We now evaluate the object composition ability of the proposed approach, PartCraft, in comparison to the baselines. In this section, we use the base generation model as a StableDiffusion 1.5 model to do a fair comparison with the Rich-Text Baseline. We use the same generation setting of using a PNDM scheduler with 41 steps and suggested classifier guidance of 8.5. We keep all the parameters the same as that of the Rich-Text model to perform a fair comparison.
Baselines. For the task of generating the object image based on part-level details of the specified object, we use the strong baselines of Rich-Text Generation and InstructPix2Pix. The Rich-Text Generation method generates the base image itself, whereas the InstructPix2Pix method requires the base image to be generated by us. Apart from this, we also evaluate the standard StableDiffusion baseline in which we add all the part details in the text prompt itself to generate the desired image. We compare all these baselines with our proposed method, PartCraft, while ensuring that the base parameters and seed are same for the base image generation. This is done to ensure a fair comparison with baselines, as random seeds often change the generation completely. We defer specific details for the baselines to the Suppl. Section.
Visual Comparison. We provide a visual comparison for the baselines in the case of the a) bird image distinctive part generation and b) human image part generations. Across both cases (Fig. 5), our proposed method, PartCraft, only modifies the desired region of the beak and replaces them with the iconic pelican beak. In comparison, the StableDiffusion baseline modifies the image completely. In the Rich-Text generation, the full region corresponding to the bird gets modified instead of only the specified prompt in base generation. For the editing-based InstructPix2Pix method, we observe that it modifies the base image at a global level and is unable to localize modification to the desired part region (Fig. 5). In general, we observe that PartCraft generated images are novel compositions, that look aesthetic and also follow the part-level instructions.
Quantitative Evaluation. As part-specific generation is fine-grained we perform both automatic and user-study evaluation. We performed a user study by inviting 50 participants (28 responded with a full survey). We evaluate the model by asking questions and evaluating metrics on three orthogonal aspects i.e., a) Localization of part generation in comparison to the base image (through LPIPS [42]), b) Text-Consistency of generated parts by measuring CLIP similarity of the part detailed text prompt with the generated image and c) ) Aesthetic Quality of the generated image through LAION-5B [30] CLIP Aesthetic scorer. In total, we collect about 3.5k opinions. In Fig. 7, we summarize the results in which we observe that PartCraft is significantly preferred over the other baselines in terms of localized and consistent Part Generation. The automated evaluation results are provided in the table below which also follow a similar trend as in the user-study, demonstrating effectiveness of PartCraft in localized part generation, while ensuring aesthetics score is similar to that of base model.
5 Analysis and Discussion
Part Diffusion Based Localization. We ablate the components we have introduced in the PartDiffusion process of PartCraft (Sec. 3.2). We provide an analysis of the effect of using a) null-text inversion , d) max-based localization c) usage of dot product-based protocol in the part assignment, and d) independent text embeddings. We have used a subset of CUB-200 images to perform all evaluations, which is kept fixed across ablations. We tabulate the ablations in Table 7. We observe that all components introduced in PartCraft significantly contribute to the overall segmentation performance of the part localization.
Generalization of PartCraft. In Fig. 8, we first have a painting of Claude Monet which is generated by the base StableDiffusion method. We then use PartCraft to specify the dress part of women’s token in the prompt, to dark green color in Monet Style, orange color dress in Van Gogh Style (middle), and green colored dress in Ukiyo-e style. In Fig. 8, we provide edits for the ‘Woody’ character from Toy Story. Across these results, we find that PartCraft is able to generate variations of shirt, pant, and dress colors and is able to generate aesthetic combinations. This zero-shot generalization across domains demonstrates the potential of creative activities that can be enabled with PartCraft.
Limitations: We find that the generation of part is bottlenecked with the part understanding of StableDiffusion. We find that if the part is localized correctly, the text-2-image model is able to generate specified details in PartCraft. Hence, improving part-level localization of these models is a good direction for future.
6 Conclusion
In this work, we introduce PartCraft a method to generate object images with fine-grained details specified at the object part level, using an expressive Rich-Text interface. The PartCraft method introduces a novel part diffusion process, which is responsible for denoising objects using the part features, and then utilizes a region-specific diffusion process to generate part details and compose the final image. PartCraft enables the training-free part-level control for models.
Supplementary Material: Crafting Parts for Expressive Object Composition
Harsh Rangwani Aishwarya Agarwal Kuldeep Kulkarni R. Venkatesh Babu Srikrishna Karanam
Appendix 0.A Notations
We provide notations used in the paper in Table 0.A.1.
Table 0.A.1: Notation Table
@@ Appendix 0.A Notations
Appendix 0.B Potential Negative Impact
Text-2-image generative models have shown great promise in image generation and can be utilized in content and media creation. However, ensuring the created content is unbiased, harmless, and free of misinformation is essential. Our work gives users more control over text-2-image generation, which should be used responsibly to avoid misinformation.
Appendix 0.C Implementation Details
We discuss implementing the two parts of the PartCraft Process, the Part Localization and Part Generation processes.
Part Localization. In Part Localization, we first discuss the implementation on StableDiffusion 2.1, where we use 50 steps of DDIM Scheduler [31] for denoising. As we aim to evaluate our method on CUB-200 and DeepFashion datasets, we first perform inversion using Null-Text Inversion [26] with a guidance scale of 0.05 (other hyper-parameters are kept default). We keep the $\alpha$ value linearly scaled with time steps from 0 to 0.5. To obtain the mask for each segment, we keep a low $\epsilon=0.05$ threshold, as here, we want to segment the region in all possible parts as the DeepFashion and CUB-200 generate localization masks for the 14 part regions. However, for comparison with baselines producing segments in 4 parts, we perform the following clustering of parts in the four cluster regions (Table 0.C.1). We use (K=9) to cluster the different regions into 4 clusters. We use the foreground masks for objects provided by Choudhury et al. [11] [^2] https://github.com/subhc/unsup-parts for evaluating the FG-NMI and FG-ARI metrics in Table 1.
Table 0.C.1: Part Names in Clusters for CUB and DeepFashion Datasets
@@ Appendix 0.C Implementation Details
We use StableDiffusion 1.5 to compare the generation results to the Rich-Text [14] [^3] https://github.com/SongweiGe/rich-text-to-image baseline. In this case, we want to start the localization process at a later stage of denoising, as we aimed the part diffusion not to alter the base generation too much; hence, start part denoising from step $T_{th}=24$. We keep $\epsilon=0.5$ as the threshold and $\delta=0.3$ for max part localization (in Eq. 6 and 4) respectively. We use a PNDM scheduler with 41 steps and an 8.5 guidance scale, as done by default in Rich-Text [14] generation.
@@
Table 0.C.2: Ablation Analysis of PartCraft for Part Localization.
PartCraft (Ablations) Method FG-NMI FG-ARI PartCraft 35.4 11.0 w/o Null-Text Inversion 23.1 5.2 w/o Max Localization 21.3 2.8 w/o Dot Product Localization 23.7 5.0 w/o Independent Text 31.2 8.9 PartCraft (Clustering) PartCraft (K = 9) 35.4 11.0 PartCraft (K = 4) 20.4 2.8 PartCraft (K = 14) 35.2 10.3
Figure 0.C.1: Normalized Attention Maps for parts which are localized (left) and non-localized (right).
@@
Baselines. For the Rich-Text baseline, we use the attribute properties $\mathbf{a}_{i}$ for the part token and add it to the base token if the part token is present in the base prompt. In other cases, we add the Part information as the footnote of the object token in the base prompt. For the InstructPix2Pix [7] baseline, we individually add the instructions to each part on the base prompt generation through stable diffusion. For the StableDiffusion baseline, we add the PartCraft instructions as the object’s description in the prompt. We use the same StableDiffusion model with the same seed and guidance scale across baselines. We run it on the same Nvidia A100 40GB to ensure the sanity of comparison across methods.
Appendix 0.D Analysis of PartCraft
@@
In this section, we provide additional analysis regarding the design choices made in PartCraft. In particular, we ablate the components we have introduced in the PartDiffusion process of PartCraft (Sec. 3.2). We provide an analysis of the effect of using a) independent text embeddings, b) usage of dot product-based protocol in the part assignment, and c) the visualization of attention maps for the localized and unlocalized parts. We have used a subset of CUB-200 images to perform all evaluations, which is kept fixed across ablations. We tabulate the ablations in Table 0.C.2 (also including ablations from the main text Table 2) and provide visual results in Fig. 0.C.1. We observe that all components introduced in PartCraft significantly contribute to the overall segmentation performance of the part localization module. Further, in Fig. 0.C.1, we find that if we normalize parts that do not satisfy localization conditions (forehead and back), it leads to high attention values in most regions. This demonstrates the effectiveness of the max-value based selection (Eq. 4) of parts proposed in PartCraft.
@@
Figure 0.D.1: Qualitative Comparison of part masks (on DeepFashion) generated from the Stable Diffusion Baseline to PartCraft (Ours). The Stable Diffusion Baseline assigns arbitrary part masks to segments, whereas our part marks are consistent to part.
@@
For comparison to the Stable Diffusion baseline, in addition to the results provided in Table 1, we provide a qualitative comparison in Fig. 0.D.1 and 0.D.2. The Stable Diffusion baseline assigns arbitrary masks to the wrong parts. On the contrary, if PartCraft localizes parts, they are often correctly associated with the right region of the object. This shows the advantage of using part diffusion rather than inducing additional tokens in the base prompt of StableDiffusion.
@@
Figure 0.D.2: Qualitative Comparison of part masks (on CUB200) generated from the Stable Diffusion Baseline to PartCraft (Ours). The Stable Diffusion Baseline assigns arbitrary part masks to segments, whereas our part marks are consistent to part.
@@
Figure 0.D.3: PartCraft is robust when the masks don’t exactly fit the specified parts. This is because the text description for the masked region contains a description only of specified parts (e.g. beak and head above). Hence, the other regions, despite being in the masked region, remain similar to the base generation.
Figure 0.D.4: Comparison of PartCraft to Other Approaches, for generating using the prompts specified in Fig. 1 of main draft. The PartCraft Approach can modify base generations in the parts specified and can generate aesthetic images in comparison to the state-of-the-art baselines.
Figure 0.D.5: A limiting case when only color is specified in PartCraft, and the masks cover extraneous regions. In such cases, when only color guidance is applied, it might leak to some other parts instead of the specified part (‘shirt’). As color is added through gradient guidance in the entire masked region.
Appendix 0.E Qualitative Analysis of PartCraft Generations
@@
Robustnes of PartCraft w.r.t. Masks. PartCraft uses masked diffusion process to generate and compose the object parts. Hence, we first see the effect of the mask region, where we observe (Fig. 0.D.3) that in cases where the mask occupies more region than the desired part, the part diffusion process mostly modifies the requested part. This demonstrates that PartCraft can still generate desired aesthetic outputs in case the masks are not very accurate.
Comparison of PartCraft to Baselines (Fig. 1). In Fig. 0.D.4, we demonstrate a comparison of the StableDiffusion, InstructPix2Pix, and Rich-Text on the prompts demonstrated in the teaser figure. For Rich-Text, we add the description in PartText to the part token if it’s present in the base prompt; if not, we add the part description as a footnote in the object token in the base prompt. For InstructPix2Pix, we make instructions regarding part modification one by one on the base generation. We find that the baselines significantly change the entire composition of the object instead of just specified parts.
@@
Limitations. We find that the color guidance applies to the entire masked region. Hence, in some cases, the font color specification $\mathbf{a}_{c}$, can alter other areas in minor ways besides the specified part. This is due to color gradient loss, also used in Rich-Text [14] to specify the specific color of the region. We highlight that in the example in Fig. 0.D.5, where some part of the orange lines appear in the shorts worn by the child as it was also the part of mask.
Acknowledgments: Harsh Rangwani is supported by the PMRF Fellowship at Indian Institute of Science. A considerable part of this work was carried out when Harsh Rangwani was an intern at Adobe Research.

### v2

@@ Appendix C Attribute Variations
Appendix D Notations
We provide notations used in the paper in Table 3.
Table 3: Notation Table for the paper.
@@ Appendix D Notations
Appendix E Potential Negative Impact
Text-to-image generative models have shown great promise in image generation and can be utilized in content and media creation. However, ensuring the created content is unbiased, harmless, and free of misinformation is essential. Our work gives users more control over text-to-image generation, which should be used responsibly to avoid misinformation.
Appendix F Implementation Details
We discuss implementing the two parts of the PartComposer process, the Part Localization and Part Generation steps.
Part Localization. In Part Localization, we first discuss the implementation on StableDiffusion 2.1, where we use 50 steps of DDIM Scheduler [43] for denoising. As we aim to evaluate our method on CUB-200 and DeepFashion datasets, we first perform inversion using Null-Text Inversion [32] with a guidance scale of 0.05 (other hyper-parameters are kept default). We keep the $\alpha$ value linearly scaled with time steps from 0 to 0.5 (Eq. 1). To obtain the mask for each segment, we keep a low $\epsilon=0.05$ threshold (Eq. 3), as here, we want to segment the region in all possible parts as the DeepFashion and CUB-200 generate localization masks for the 14 part regions. However, for comparison with baselines producing segments in 4 parts, we perform the following clustering of parts in the four cluster regions (Table 4). We use (K=9) to cluster the different regions into 4 clusters. We use the foreground masks for objects provided by Choudhury et al. [12] [^3] https://github.com/subhc/unsup-parts for evaluating the FG-NMI and FG-ARI metrics in Table 4.
Table 4: Part Names in Clusters for CUB and DeepFashion Datasets
@@ Appendix F Implementation Details
Table 5: Ablation Analysis of PartComposer for Part Localization.
PartComposer (Ablations) Method FG-NMI FG-ARI PartComposer 35.4 11.0 w/o Null-Text Inversion 23.1 5.2 w/o Max Localization 21.3 2.8 w/o Dot Product Localization 23.7 5.0 w/o Independent Text 31.2 8.9 PartComposer (Clustering) PartComposer (K = 9) 35.4 11.0 PartComposer (K = 4) 20.4 2.8 PartComposer (K = 14) 35.2 10.3
Figure 14: Normalized Attention Maps for parts which are localized (left) and non-localized (right).
We use StableDiffusion (SD) 1.5 to compare the generation results to the Rich-Text [18] [^4] https://github.com/SongweiGe/rich-text-to-image baseline. In this case, we want to start the localization process at a later stage of denoising, as we aimed the part diffusion not to alter the base generation too much; hence, start part denoising from step $T_{th}=24$. We do merging of base with initial diffusion process for $0.5$ number of steps (Eq. 5) and $\delta=0.3$ for max part localization (in Eq. 2) respectively. We use a PNDM scheduler with 41 steps and an 8.5 guidance scale, as done by default in Rich-Text [18] generation.
@@
Baselines. For the Rich-Text baseline, we use the attribute properties $\mathbf{a}_{i}$ for the part token and add it to the base token if it is in the base prompt. In other cases, we add the Part information as the footnote of the object token in the base prompt. For the InstructPix2Pix [8] baseline, we iteratively add the instructions for each part on the base generation through. For the StableDiffusion baseline, we add the PartComposer instructions as the object’s description in the prompt. We use the same StableDiffusion model with the same seed and guidance scale across baselines. We run it on the same Nvidia A100 40GB to ensure the sanity of comparison across methods. PartCraft [33], a recent method, also aims to generate objects based on part composition. However, if the text-to-image model is PartCraft trained for CUB-200, it can only generate variations of CUB-200. Further, it also requires annotations of the parts regarding key points, etc. Hence, it cannot generate parts based on text descriptions like our work. Hence, due to additional supervision and focus on particular datasets, it cannot be used as a general baseline for comparison with our work.
Appendix G Analysis of PartComposer
@@
In this section, we provide additional analysis regarding the design choices made in PartComposer. In particular, we ablate the components we have introduced in the PartDiffusion process of PartComposer (Sec. 3.2). We provide an analysis of the effect of using a) independent text embeddings, b) usage of dot product-based protocol in the part assignment, and c) the visualization of attention maps for the localized and unlocalized parts. We have used a subset of CUB-200 images to perform all evaluations, which is kept fixed across ablations. We tabulate the ablations in Table 5 (also including ablations from the main text Table 2) and provide visual results in Fig. 14. We observe that all components introduced in PartComposer significantly contribute to the overall segmentation performance of the part localization module. Further, in Fig. 14, we find that normalizing parts that do not satisfy localization conditions (forehead and back) leads to high attention values in most regions. This demonstrates the effectiveness of the max-value-based selection (Eq. 4) of parts proposed in PartComposer.
@@
Figure 15: Qualitative Comparison of part masks (on DeepFashion) generated from the Stable Diffusion Baseline to PartComposer (Ours). The Stable Diffusion Baseline assigns arbitrary part masks to segments, whereas our part marks are consistent to part.
@@
For comparison to the Stable Diffusion baseline, in addition to the results provided in Table 4, we provide a qualitative comparison in Fig. 15 and 16. The Stable Diffusion baseline assigns arbitrary masks to the wrong parts. On the contrary, if PartComposer localizes parts, they are often correctly associated with the right region of the object. This shows the advantage of using part diffusion rather than inducing additional tokens in the base prompt of StableDiffusion.
@@
Figure 16: Qualitative Comparison of part masks (on CUB200) generated from the Stable Diffusion Baseline to PartComposer (Ours). The Stable Diffusion Baseline assigns arbitrary part masks to segments, whereas our part marks are consistent to part.
@@
Figure 17: PartComposer is robust when the masks don’t exactly fit the specified parts. This is because the text description for the masked region contains a description only of specified parts (e.g. beak and head above). Hence, the other regions, despite being in the masked region, remain similar to the base generation.
@@
Robustnes of PartComposer w.r.t. Masks. PartComposer uses masked diffusion process to generate and compose the object parts. Hence, we first see the effect of the mask region, where we observe (Fig. 17) that in cases where the mask occupies more region than the desired part, the part diffusion process mostly modifies the requested part. This demonstrates that PartComposer can still generate desired aesthetic outputs in case the masks are not very accurate.
Appendix H Limitations of PartComposer Generations
Figure 18: Limiting Case. When only color is specified in PartComposer, and the masks cover extraneous regions. In such cases, when only color guidance is applied, it might leak to some other parts instead of the specified part (‘shirt’). As color is added through gradient guidance in the entire masked region.
@@ Appendix H Limitations of PartComposer Generations
We find that the color guidance applies to the entire masked region. Hence, in some cases, the font color specification $\mathbf{a}_{c}$, can alter other areas in minor ways besides the specified part. This is due to color gradient loss, also used in Rich-Text [18] to specify the specific color of the region. We highlight that in the example in Fig. 18, where some part of the orange lines appear in the shorts worn by the child as it was also the part of the mask. We find that mask localization is the bottleneck for performance. Hence, improving the part understanding and segmentation capability of the Text-to-Image models is an important direction to be pursued across future works, including explorations on new models like Flux and SD3.5.